OpenStudy (anonymous):

Melissa is 24 years younger than Joyce. In 2 years, Joyce will be 3 times as old as Melissa will be then. How old are they now?

OpenStudy (anonymous):

I'm working on it

OpenStudy (callisto):

Let x be the age of Melissa now. From the question, ''Melissa is 24 years younger than Joyce''. That means age of Joyce is age of Melissa + 24, so age of Joyce = x+24. Two years later... Melissa will be x+2 years old. Joyce will be (x+24) + 2 years old. By then, Joyce will be 3 times as old as Melissa, So, 3 times the age of Melissa two years later will be the age of Joyce two years later. So, you get the equation 3 (x+2) = (x+24) + 2 And you have to solve it. Do you understand what I've been working on?

OpenStudy (anonymous):

3x + 6 = x +24 +2 3x - x = 26-6 2x = 20 x = 10 But Melissa is x +2 years old Melissa is 10 +2 years old Therefore, Melissa is 12 years old

OpenStudy (anonymous):

Joyce = (x +24) +2 years old Joyce = 10 + 24 +2 Joyce = 36 years old Therefore, Melissa is 12 Years old while Joyce is 36 years old. The answer is accurate with the question. Given that Melissa is 24 years younger